### v2

# Verification and Refinement of Natural Language Explanations through LLM-Symbolic Theorem Proving

Abstract
Natural language explanations have become a proxy for evaluating explainable and multi-step Natural Language Inference (NLI) models. However, assessing the validity of explanations for NLI is challenging as it typically involves the crowd-sourcing of apposite datasets, a process that is time-consuming and prone to logical errors. To address existing limitations, this paper investigates the verification and refinement of natural language explanations through the integration of Large Language Models (LLMs) and Theorem Provers (TPs). Specifically, we present a neuro-symbolic framework, named Explanation-Refiner, that augments a TP with LLMs to generate and formalise explanatory sentences and suggest potential inference strategies for NLI. In turn, the TP is employed to provide formal guarantees on the logical validity of the explanations and to generate feedback for subsequent improvements. We demonstrate how Explanation-Refiner can be jointly used to evaluate explanatory reasoning, autoformalisation, and error correction mechanisms of state-of-the-art LLMs as well as to automatically enhance the quality of human-annotated explanations of variable complexity in different domains.
1 Introduction
A recent line of research in Natural Language Inference (NLI) focuses on developing models capable of generating natural language explanations in support of their predictions (Thayaparan et al., 2021; Chen et al., 2021; Valentino et al., 2022; Bostrom et al., 2022; Weir et al., 2023). Since natural language explanations can be used as a proxy to evaluate the underlying reasoning process of NLI models (Kumar and Talukdar, 2020; Zhao and Vydiswaran, 2021; Chen et al., 2021), researchers have proposed different methods for assessing their intrinsic quality (Wiegreffe and Marasovic, 2021; Camburu et al., 2020; Valentino et al., 2021; Atanasova et al., 2023; Quan et al., 2024; Dalal et al., 2024), including the adoption of language generation metrics for a direct comparison between models’ generated explanations and human-annotated explanations.
However, this process is subject to different types of limitations. First, the use of language generation metrics requires the crowd-sourcing of explanation corpora to augment existing NLI datasets Wiegreffe and Marasovic (2021), a process that is typically time-consuming and susceptible to errors (Liu et al., 2022; Zhao et al., 2023). Second, language generation metrics have been shown to fail capturing fine-grained properties that are fundamental for NLI such as logical reasoning, faithfulness, and robustness (Atanasova et al., 2023; Camburu et al., 2020; Chan et al., 2022; Quan et al., 2024). Third, human explanations in NLI datasets tend to be incomplete and contain logical errors that could heavily bias the evaluation Elazar et al. (2021); Valentino et al. (2021).
In this paper, we investigate the integration of state-of-the-art LLM-based explanation generation models for NLI with external logical solvers to jointly evaluate explanatory reasoning Pan et al. (2023a); Olausson et al. (2023); Jiang et al. (2024b) and enhance the quality of crowd-sourced explanations. In particular, we present a neuro-symbolic framework, named Explanation-Refiner, that augments a Theorem Prover (TP) with Large Language Models (LLMs) to investigate the following research questions: RQ1: “Can the integration of LLMs and TPs provide a mechanism for automatic verification and refinement of natural language explanations?”; RQ2: “Can the integration of LLMs and TPs improve the logical validity of human-annotated explanations?”; RQ3: “To what extent are state-of-the-art LLMs capable of explanatory reasoning, autoformalisation, and error correction for NLI in different domains?”. To answer these questions, Explanation-Refiner employs LLMs to generate and formalise explanatory sentences and to suggest potential inference strategies for building non-redundant, complete, and logically valid explanations for NLI. In turn, the TP is adopted to verify the validity of the explanations through the construction of deductive proofs and the generation of fine-grained feedback for LLMs.
We instantiate Explanation-Refiner with state-of-the-art LLMs (i.e., GPT-4 (OpenAI, 2023), GPT-3.5 (Brown et al., 2020), LLama (Touvron et al., 2023), and Mistral (Jiang et al., 2024a)) and the Isabelle/HOL proof assistant (Nipkow et al., 2002) utilising Neo-Davidsonian event semantics (Parsons, 1990) coupled with First-Order Logic (FOL) to effectively and systematically translate natural language sentences into logical forms.
Our empirical analysis carried out on three NLI datasets of variable complexity (i.e., e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018)) reveals that external feedback from TPs is effective in improving the quality of natural language explanations, leading to an increase in logical validity using GPT-4 from 36% to 84%, 12% to 55%, and 2% to 37% (on e-SNLI, QASC, and WorldTree respectively). At the same time, the results demonstrate that integrating external TPs with LLMs can reduce errors in autoformalisation, with an average reduction of syntax errors of 68.67%, 62.31%, and 55.17%. Finally, we found notable differences in performance across LLMs and NLI datasets, with closed-sourced LLMs (i.e., GPT-4 and GPT-3.5) significantly outperforming open-source models (i.e., Mistral and LLama) on both explanatory reasoning and autoformalisation, along with a shared tendency of LLMs to struggle with increasing explanation complexity.
Figure 1: The overall pipeline of Explanation-Refiner: An NLI problem is converted into axioms and theorems as a theorem prover’s theory, along with some proof steps derived from a preliminary inference to send to a theorem prover. In case the proof fails (logically invalid), the erroneous step, along with the proof strategy and proof steps, are extracted as feedback to refine the explanatory sentences in a new iteration.
To summarise, the main contributions of this paper are:
  1. 1.
    We introduce Explanation-Refiner, a novel neuro-symbolic framework that integrates LLMs with an external theorem prover. This framework automatically verifies and refines explanatory sentences in NLI tasks using an objective external feedback.
  2. 2.
    We utilise Neo-Davidsonian event semantics coupled with FOL to effectively translate natural language sentences into logical forms to minimises semantic information loss. Additionally, we introduce a novel method that leverages a theorem prover and a proof assistant for verifying NLI explanations and a syntactic refiner to minimise syntax errors in responses generated by LLMs.
  3. 3.
    We conduct a comprehensive series of experiments to assess Explanation-Refiner across five LLMs on three datasets, which include between 1 to 16 explanatory sentences. These experiments span a range of tasks from simple textual entailment to complex multiple-choice question answering in different context domains.
  4. 4.
    We perform extensive quantitative and qualitative analyses to explore the explanation refinement process. Our investigations delve into the LLMs’ inference capabilities, revealing the strengths and limitations of different models in producing verifiable and explainable logical reasoning for NLI.
2 Explanation Verification and Refinement
Explanation-based NLI is widely adopted to evaluate the reasoning process of multi-step inference models via the construction of natural language explanations. In this work, we refer to the following formalisation for Explanation-based NLI: given a premise sentence $p_{i}$, a hypothesis sentence $h_{i}$, and an explanation $E_{i}$ consisting of a set of facts $\{f_{1},f_{2},...,f_{n}\}$, the explanation $E_{i}$ is logically valid if and only if the entailment $p_{i}\cup E_{i}\models h_{i}$ holds. This entailment is considered verifiable if $\{p_{i},E_{i},h_{i}\}$ can be translated into a set of logical forms $\Phi$ that compose a theory $\Theta$. The validity of this theory, $\Theta$, is subsequently determined by a theorem prover, verifying whether $\Theta\vDash\psi$, where $\psi$ represents a logical consequence derived from the logical form of $h_{i}$.
We aim to automatically verify the logical validity of explanation $E_{i}$, and if $\Theta\vDash\psi$ is rejected by the theorem prover, a further refinement stage should be initiated to refine the facts $\{f_{1},f_{2},...,f_{n}\}$ based on external feedback, resulting in an updated explanation $E_{i}^{\prime}$. Thus, an explanation is accepted if all the facts are logically consistent, complementary and non-redundant to support the derivation.
3 Explanation-Refiner
To verify the logical validity and refine any logical errors in explanatory sentences for NLI tasks, we present a neuro-symbolic framework, Explanation-Refiner, to iteratively check and refine the explanation $E_{i}$ based on external feedback; Isabelle/HOL (Nipkow et al., 2002) proof assistant is applied as the external tool to make the logical deduction and provide feedback for explanation refinement.
Figure 1 shows an overview of our proposed framework. Given an NLI task, to evaluate the logical validity of the entailment, the LLM is prompted to perform an autoformalisation process that transforms natural language sentences into formal language represented in the form of Isabelle/HOL theory. Each fact $f\in E_{i}$ is converted into an axiom $a_{i}$, where each $a_{i}$ is an element of the set $A=\{a_{1},a_{2},...,a_{n}\}$. The premise $p_{i}$ and corresponding hypothesis $h_{i}$, is converted into a theorem for proving $p_{i}\wedge B\rightarrow h_{i}$, where $B\subseteq A$. A syntax refinement mechanism is subsequently applied to the previously transferred symbolic forms. The theorem prover is implemented as a checker to identify any syntax errors and provide these error details as feedback to an LLM, enabling the LLM to iteratively correct the syntax errors over a fixed number of iterations, denoted by $t$.
With the proof assistant, we can perform automated reasoning in the theorem prover by constructing proof steps. Step 3 initially generate a rough inference to state a preliminary proof strategy in natural language sentences and elicit the facts $f\in E_{i}$ which are non-redundant and essential for entailing the hypothesis $h_{i}$. Based on this preliminary proof strategy, the LLM is prompted to construct and formalise the proof steps for proving the theorem.
In step 5, the theorem prover will verify the constructed theory by attempting to prove the theorem. If it is solvable, we state it as a logically valid explanation. If the prover failed at one of the proof steps, we extract this failed step along with the applied axioms $B\subseteq A$ as the external feedback for an LLM. This feedback is used to refine the logical error within B and consequently refine the facts $f\in E_{i}$, which were previously converted from natural language.
theorem hypothesis:
(* Premise: A smiling woman is playing the violin in front of a turquoise background. *)
assumes asm: "Woman x $\wedge$ Violin y $\wedge$ Background z $\wedge$ Turquoise z $\wedge$ Smiling x $\wedge$ Playing e $\wedge$ Agent e x $\wedge$ Patient e y $\wedge$ InFrontOf x z"
(* Hypothesis: A woman is playing an instrument. *)
shows $\exists$ x y e. Woman x $\wedge$ Instrument y $\wedge$ Playing e $\wedge$ Agent e x $\wedge$ Patient e y"
Figure 2: An example of representing the premise and hypothesis sentences in Isabelle/HOL theorem
proof -
from asm have "Woman x $\wedge$ Violin y $\wedge$ Playing e $\wedge$ Agent e x $\wedge$ Patient e y" by blast
then have "Woman x $\wedge$ Instrument y $\wedge$ Playing e $\wedge$ Agent e x $\wedge$ Patient e y" using explanation_1 by blast
then show ?thesis using asm by blast
qed
Figure 3: An example of proof constructed by the LLM for the Isabelle/HOL proof assistant to verify the hypothesis.
3.1 Autoformalisation
In order to formally verify the logical validity of the explanations, we adopted Neo-Davidsonian event-based semantics and FOL.
Neo-Davidsonian Event Semantics
Preventing the loss of semantic information during the representation of natural language sentences in logical forms, such as FOL, poses significant challenges when using LLMs, particularly with long and complex sentences that are crucial for logical reasoning (Olausson et al., 2023). Neo-Davidsonian event semantics (Parsons, 1990) utilises event variables to represent the verb predicates and their corresponding object arguments as semantic roles. This approach establishes a predicate-argument structure that preserves the information content and faithfulness of complex sentences, closer to the surface form of the sentence. For example, the sentence ‘A wolf eating a sheep is an example of a predator hunting prey’ can be formalised as follows:
$\forall xye_{1}(\text{wolf}(x)\wedge\text{sheep}(y)\wedge\text{eating}(e_{1})$ (1)
$\quad\wedge\text{agent}(e_{1},x)\wedge\text{patient}(e_{1},y)\rightarrow$
$\quad(\exists e_{2}\;\text{predator}(x)\wedge\text{prey}(y)\wedge$
$\quad\text{hunting}(e_{2})\wedge\text{agent}(e_{2},x)\wedge$
$\quad\text{patient}(e_{2},y)\wedge\text{example}(e_{1},e_{2})))$
In 1, the verbs are represented as the events ‘eating’ and ‘hunting,’ where the agent and patient arguments correspond to the entities performing and receiving the actions within these events, respectively. The logical form $\text{example}(e_{1},e_{2})$ explicitly captures the semantic meaning of this sentence: the event of a wolf eating a sheep as an exemplar of a predator hunting prey. Similarly, whenever there are no action verbs involved in a sentence, we utilise FOL to represent the static or descriptive aspects. For instance:
$\forall x(\text{gravity}(x)\rightarrow\text{force}(x))$ (2)
$\forall xy(\text{greater}(x,y)\rightarrow\text{larger}(x,y))$ (3)
The above logical forms correspond to the sentences ‘gravity is a kind of force’ and ‘greater means larger’, respectively.
Isabelle/HOL Theory Construction
For the Isabelle/HOL theorem prover, a theory script is essential to facilitate the proof of a theorem. Therefore, we designate explanatory sentences as axioms:
(* Explanation 1: A violin is an instrument. *)
axiomatization where
[…]
Additionally, as illustrated in Figure 2, both the premises and hypothesis are defined as constituting parts of the theorem. The ‘assumes asm’ clause generally comprises unquantified, specific propositions or conjunctions of propositions, which are recognised as known truths. The ‘show’ clause denotes the conclusion (hypothesis) for which we seek to establish proof through logical deductions based on the assumed propositions and axioms.
Syntax Error Refiner
Recent studies (Gou et al., 2024; Olausson et al., 2023) have revealed persistent syntax errors when prompting LLMs for code and symbolic form generation tasks. The method proposed by Pan et al. (2023a) utilises error messages from symbolic solvers to iteratively refine LLM outputs. Following a similar approach, we categorised the syntax errors into two distinct subdomains based on feedback from Isabelle: type unification errors and other syntax errors. Type unification errors primarily arise from mismatches between declared and actual argument types in logical clauses. Other syntax errors typically involve missing brackets, undefined entity names, or invalid logical symbols. Our process involves using Isabelle to identify syntax errors in the transferred theory, extracting these error messages, and then prompting the LLM with these messages along with few-shot examples. This guides the model on how to correct each type of syntax error over a series of iterations, allowing for continuous verification and refinement. Details of the autoformalisation prompts are described in Appendix A.4.1.
3.2 Proof Construction
A proof provides a detailed, step-by-step strategy that elucidates the logical connections and unification among axioms to support the reasoning process aimed at achieving the solver’s goal. Initially, we prompt the LLM to create a preliminary proof in natural language to assess how it infers the hypothesis and to identify which explanatory sentences are relevant, redundant, or unrelated. Based on this initial proof, we then guide the LLM to develop an Isabelle/HOL proof (figure 3) that utilise the Isabelle/HOL proof assistant to clearly demonstrate the explanatory sentences (axioms) required to prove the hypothesis. The general proof steps generated by an LLM are in the format ‘show $X$ using $Y$ by $Z$’, where the theorem prover is asked to prove $X$ given the assumptions $Y$, using the automated proof tactic $Z$. The proof tactic often applied is ‘blast’, which is one of Isabelle’s powerful FOL theorem proving tactics, enabling efficient and automated proof discovery across a range of logical forms (Paulson, 1999). Details of proof construction process prompts are described in Appendix A.4.2.
3.3 Verify and Refine
Finally, the constructed theory, which includes axioms, theorems, and proof steps, is submitted to the theorem prover for verification. If the theory is validated, it outputs a logically sound explanation. If the proof fails or timeouts, we extract the first error from the solver’s error message, identify the corresponding proof step, and locate the related explanatory sentences (axioms) from the theory. We begin by removing redundant and irrelevant facts that are not present in the preceding Isabelle/HOL proof steps or are declared as such in the text inference strategy. Then, we prompt the LLM to refine the explanatory sentences by providing it with the error message, the failed proof step, the associated proof strategy, and the relevant explanatory sentences for further iteration. This process is iterative and progressive; with each iteration, the framework addresses one or more logical errors, continually refining the explanatory sentences to ultimately yield a logically valid and verifiable explanation. Additional details on the prompts used for refinement are described in Appendix A.4.3.
4 Empirical Evaluation
Figure 4: The initial and final number of logically valid explanations, along with the average iteration times required to refine an explanation for each LLM
Figure 5: Number of successfully refined explanations at each iteration step.
Figure 6: The average number of theories containing syntactic errors before and after the syntax refinement process
Figure 7: AF represents the autoformalisation components, and TI represents the textual inference components. TI+AF (Base Model) indicates the use of the base model for both the autoformalisation and textual inference components. TI+AF (GPT-4) indicates the use of GPT-4 for the autoformalisation components, while the base model is used for textual inference.
Figure 8: Average of proof steps processed by the proof assistant against the total proof steps suggested by the LLMs in refined and unrefined explanations.
To assess the effectiveness of integrating LLMs with external feedback for refinement purposes, we evaluated the Explanation-Refiner on two NLI tasks: textual entailment and multiple-choice question answering. For textual entailment, given a premise $p_{i}$ and an explanatory sentence as explanation $E_{i}$, the goal is to determine whether $p_{i}$ and $E_{i}$ together entail a hypothesis $h_{i}$. In the multiple-choice question answering task, there is a question $q$ accompanied by a set of candidate answers $C=\{c_{1},c_{2},...,c_{n}\}$, with $c_{i}$ identified as the correct answer. We convert $q$ and the correct answer $c_{i}$ as the hypothesis $h_{i}$. Explanatory facts serve as evidence supporting $c_{i}$ as the correct answer and are denoted as $E_{i}$, with the question’s context sentence as the premise $p_{i}$.
4.1 Datasets
We adopted three different NLI datasets for evaluation: e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018), using a total of 300 samples selected via the sampling strategy defined in (Valentino et al., 2021), which maximises representativeness and mutual exclusivity across syntactic and semantic features expressed in the datasets. e-SNLI is a crowd-sourced dataset typically used as a benchmark for textual entailment. It comprises one premise sentence, one explanatory sentence, and one hypothesis sentence per sample. QASC and WorldTree, on the other hand, are datasets designed for multiple-choice question answering within the scientific domain. QASC includes two explanatory sentences for each correct answer. WorldTree is the most complex of the three, featuring between 1 and 16 explanatory sentences in each sample.
4.2 Theorem Prover
We adopted Isabelle/HOL (Nipkow et al., 2002) proof assistant as the theorem prover. To integrate this theorem prover as a real-time verification tool with LLMs, we employ a Python client (Shminke, 2022) as TCP (Transmission Control Protocol) client to configure Isabelle/HOL as a server. This enables the communication of the constructed theory files and the extraction of the response messages from Isabelle.
4.3 Models
We conducted experiments using five LLMs within the proposed framework. The models include two open-sourced models: Llama2-70b (Touvron et al., 2023) and Mixtral-8x7b (Jiang et al., 2024a), as well as Mistral-Small (mistral-small-latest) (Mistral AI, 2024), GPT-Turbo-3.5 (gpt-3.5-turbo) (Brown et al., 2020), and GPT-4 (gpt-4-0613) (OpenAI, 2023). Temperature settings were adjusted to 0 for GPT-Turbo-3.5 and GPT-4, and to 0.01 for Llama2-70b, Mixtral-8x7b, and Mistral-Small, aiming to achieve both determinism in the output and effective code generation for theorem prover.
4.4 Results
Providing feedback on the exact failure step from an external theorem prover effectively guides LLMs in continuously verifying and refining explanations in NLI tasks.
To assess the effectiveness of employing an external theorem prover to verify and refine explanations in NLI tasks, we conducted a comparative analysis across various LLMs (Figure 4). The initially valid explanations represent the percentage of explanations that can be verified as logically valid without any further iteration. Although the initial verification results varied among different models, all LLMs demonstrated a consistent improvement in refining the logical validity of the explanations. This process highlights the positive impact of the external feedback but also shows significant differences between models. We found that lower rates of initial valid explanations often resulted from syntactic errors, which impeded the theorem prover’s ability to generate proofs. Despite this initial variability, all models demonstrate a consistent improvement in the refinement process across the datasets. Notably, GPT-4 outperformed other models, improving the validity of explanations by 48%, 43%, and 35% across the three datasets, respectively, within a maximum number of ten iterations (Figure 4). Figure 5 shows the number of explanations refined at each iteration across the e-SNLI, QASC, and WorldTree datasets. On average, we found that an increasing number of iterations leads to increasing refinement, with models requiring an average of five iterations across the datasets.
Explanation length/complexity impacts formalisation.
The e-SNLI dataset, which includes only a single explanatory sentence per example, shows the best overall performance. In contrast, the multiple-choice question answering datasets, QASC and WorldTree, exhibit comparatively lower performance. QASC typically contains 2 explanatory sentences, while WorldTree ranges from 1 to 16 sentences. As the number of explanatory sentences increases, so does the complexity of the logical reasoning required. The WorldTree dataset, in particular, poses the greatest challenge due to its demand for multi-hop inference strategies. Models show lower refining performance in WorldTree when compared to e-SNLI and QASC, with only 3%, 5%, and 6% of Llama-70b, Mixtral-8x7b, and Mistral-Small explanations being refined in WorldTree. Meanwhile, 29% and 37% of explanations are refined by GPT-3.5-Turbo and GPT-4 in WorldTree, respectively. This process involves synthesising multiple explanatory sentences to fulfill sub-goals, which must then be integrated to meet the overall hypothesis goal.
Iterative and categorical refinement can monotonically reduce syntax errors in responses generated by LLMs.
To evaluate the syntax error refinement stage, we quantified the presence of syntax errors in the Isabelle theories both before and after the iterative refinement process. After a maximum of three iterations, all models showed significant reductions, with maximum reductions of 68.67%, 62.31%, and 55.17% from 7.82 to 2.45, 20.27 to 7.64, and 22.91 to 10.27 across the three respective datasets (see Figure 6). While models like Llama2-70b and Mixtral-8x7b still exhibit some syntax errors in the refined theories’ code, this is primarily due to their inability to perform complex autoformalisation, especially for multiple and more complex explanatory sentences such as those in the WorldTree dataset. This result is consistent with the percentage of explanations that were successfully refined across the models, which suggests that the autoformalisation process plays a critical role in the models’ logical reasoning capability.
4.5 Ablation Study
We conducted an ablation study to further evaluate and disentangle the impact of autoformalisation on performance. To this end, we adopted GPT-4 exclusively for the autoformalisation component, while retaining the original models for explanation refinement and proof strategy generation. As shown in Figure 7, integrating GPT-4 for autoformalisation led to a significant increase in the number of explanations successfully refined across all models. For instance, Llama2-70b with GPT-4 as the formalisation component refined explanations from 7% to 65% in the e-SNLI dataset. For the multiple-choice question answering dataset, GPT-3.5-Turbo showed a relatively smaller increase from 44% to 48% and from 29% to 34%. Despite these improvements, a performance gap persists between GPT-4 and the other models, which is attributed to GPT-4’s superior symbolic reasoning capabilities required for explanation refinement from the identified logical errors.
Explanations are progressively made more complete and consistent through iterative symbolic refinement.
In order to deliver step-wise logical consistency, explanations need to be made complete and self-contained, leading to the introduction of additional explanatory sentences, leading to an increase in the total number of suggested proof steps. Therefore, we further evaluated how the proof steps vary when the total number of suggested proof steps increases contrasting both refined and unrefined cases. Figure 8 illustrates this trend. In general, all models show a positive trend, as the total suggested proof steps increase, the average number of proof steps processed by the proof assistant also increases. Models like Mistral-Small and GPT-3.5-Turbo tend to suggest more proof steps to accomplish the logical goal, which can result in some redundant steps, such as the significant pulse shown in Figure 8(c). For unrefined explanations, as shown in Figure 8(d), 8(e) and 8(f), the progression is steadier but retains a positive trend, where the models generally suggest more proof steps in response to the additional explanatory sentences introduced to correct a logical error identified from the erroneous step. We also conducted experiments on the relationship of average number of successfully processed explanatory sentences in one proof against total planned explanatory sentences in a suggest proof in appendix A.3. Examples of refined and unrefined explanations can be found in Appendix A.5.
Figure 9: Human evaluation of refined explanations in terms of factuality and triviality. Factually Correct indicates the percentage of explanation sentences that are correct in terms of commonsense knowledge. Not Trivial states the percentage of explanations that contain only explanatory sentences that repeat or paraphrase the premise and/or the hypothesis.
4.6 Factual Errors and Trivial Explanations
In addition to evaluating the logical validity of explanations, we also conducted a human evaluation of the refined explanations, considering factual correctness and explanation triviality, for the two best-performing models (GPT-3.5-Turbo and GPT-4). This evaluation focused on two questions: “Are the refined explanatory sentences factually correct?” and “Is the explanation trivial, merely repeating or paraphrasing the content of the premise and hypothesis to achieve logical validity?”. As illustrated in Figure 9, our findings indicate that all refined explanations in the e-SNLI and WorldTree datasets are consistent with commonsense knowledge. In the QASC dataset, 2.27% and 1.82% of the explanation refined by GPT-3.5-Turbo and GPT-4 contain sentences misaligned with true world knowledge. We found that the majority of these errors result from over-generalisation, such as the sentence All tetrapods are defined to have four limbs, which inaccurately includes snakes.
Finally, we found a relatively low number of explanations that repeat or paraphrase the content of premise and hypothesis. This phenomenon is absent in e-SNLI and becomes more evident when the explanatory sentences increase in complexity (i.e., WorldTree), leading models sometimes to generate explanations that do not include any additional information for the entailment to hold.
5 Related Work
5.1 LLMs Self-Refinement from External Feedback
Self-refinement of LLMs has demonstrated promising effectiveness in generating faithful and trustworthy responses (Pan et al., 2023b). The use of external feedback to guide LLMs has been extensively studied (Olausson et al., 2024a; Yu et al., 2023; Akyurek et al., 2023). Previous work such as Peng et al. (2023) and Li et al. (2024) have employed facts retrieved from external knowledge bases as sources of feedback, while Paul et al. (2024) developed a critic model to provide feedback for reasoning refinement. Additionally, Nathani et al. (2023) have explored the use of feedback models for automated feedback generation. Various works have also investigated tasks related to code generation (Chen et al., 2023; Olausson et al., 2024b) and the creation of either synthetic or expert-written logical natural language expressions (Olausson et al., 2023). Quan et al. (2024) leverage a differentiable logic reasoner to verify and refine explanations through abductive reasoning , enhancing the logical consistency of explanations in ethical NLI tasks based on the solver’s output. This paper focuses on the automated refinement of natural language sentences created by human annotators. It verifies and refines the logical validity of these human-generated sentences using detailed external feedback, which can identify the exact erroneous steps to effectively refine logical errors in the explanatory sentences.
5.2 Autoformalisation
Autoformalisation refers to the process of translating natural language descriptions into symbolic representations. Research in this area has included the formalisation of mathematical proofs (Cunningham et al., 2022; Wu et al., 2022; First et al., 2023), and efforts to transform natural language sentences into logical forms using LLMs (Pan et al., 2023a; Olausson et al., 2023; Jiang et al., 2024b; Dalal et al., 2024). However, contextual information is frequently lost when sentences are translated in these logical frameworks. To mitigate semantic loss during the transformation process, we leverage Neo-Davidsonian event semantics, which aims to maximise the preservation sentence-level content. This representation paradigm can facilitate a more systematic content-preserving translation to logical forms, which is more independent from particular choices of representation schemas.
5.3 Conclusion
In this work, we present a novel neuro-symbolic framework that automatically verifies and refines natural language explanations using iterative refinement cycles between LLMs and theorem provers (TPs). We have conducted extensive experiments on both textual entailment and multiple-choice question answering tasks, demonstrating that the proposed method, Explanation-Refiner, effectively enhances the logical validity of such human-annotated explanations. We investigated the model’s performance from simple to complex explanatory/sentence structures and introduced a method to prevent the loss of semantic information in autoformalisation tasks with error correction. Furthermore, we investigated the iterative model at each iteration to evaluate how the explanation is refined at each cycle. In future work, we aspire to enhance the framework’s robustness towards complex and unstructured explanations with fewer iterations required to improve the model’s efficiency.
5.4 Limitations
While this work have demonstrated significant improvements in terms of enhancing the logical consistency of explanations, the connection between improved logical consistency and AI safety still needs further investigation. While the concept of using formal solvers in conjunction with LLMs delivers a promise avenue to improve the consistency of reasoning within LLMs, these methodologies needs to be further developed and critically assessed as a mechanism which can provide guarantees of correctness, consistency and completeness within critical application domains.
Acknowledgments
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath (200021_204617), by the EPSRC grant EP/T026995/1, “EnnCore: End-to-End Conceptual Guarding of Neural Architectures” under Security for all in an AI enabled society, by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
Appendix A Appendix
A.1 Algorithm
Algorithm 1 shows the overall framework of Explanation-Refiner.
A.2 Scalability
Figure 10 shows the average Isabelle/HOL solving time against the number of planned explanatory sentences in a proof and the length of suggested proof steps, including theories that have syntax errors, respectively. In some cases, the theorem prover may get stuck on a proof step, and we have set a termination time if the solving time exceeds 65 seconds.
Figure 10: (a) Average Isabelle/HOL solving time against number of explanatory sentences planned in a proof. (b) Average Isabelle/HOL solving time against number of suggested proof steps in a proof.
A.3 Average Processed vs. Planned Explanatory Sentences per Proof
Figure 11 and Figure 12 shows experiments on average number of successfully processed explanatory sentences in one proof against total planned explanatory sentences in a suggest proof. Figure 13 also shows the comparison of average processed proof steps against total suggested proof steps in all dataset.
A.4 Prompts
A.4.1 Autoformalisation
Figure 14 displays the prompts used to identify action verbs (events) within the premise, explanation, and hypothesis sentences, representing events in Davidsonian-event semantics. Figure 15 displays the prompts used to transfer natural language to logical forms based on the identified events verbs. Figure 16 shows how to convert logical forms into Isabelle/HOL code (axioms and type declaration). Figure 17 shows how to convert the premise and hypothesis sentences into the Isabelle/HOL theorem code, based on the previously constructed axioms code. Figure 18 shows how to refine the syntax errors based on the types of errors, the provided code, the error messages, and the locations of the errors within the code.
A.4.2 Proof Construction
Figure 19 shows the prompts for making a preliminary inference strategy, which also identifies redundant and related explanatory sentences that will be used for proof generation. Figure 20 shows the prompts for building the proof steps used for Isabelle/HOL Proof assistant based on the provided inference strategy.
A.4.3 Explanation Refinement
Figure 21 shows how to refine the explanatory sentences based on the provided information.
A.5 Examples of Explanation Refinement
A.5.1 e-SNLI Refined Examples
Table 1 shows an example from the e-SNLI dataset of how the explanation changes after each iteration. Figures 22, 23, and 24 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 2 along with Figures 25 and 26, and Table 3 with Figures 27, 28, and 29 are two more examples.
A.5.2 QASC Refined Examples
Table 4 shows an example from the QASC dataset of how the explanation changes after each iteration. Figures 30, 31 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 5 along with Figures 32 and 33, and Table 6 with Figures 34, 35 are two more examples.
A.5.3 WorldTree Refined Examples
Table 7 shows an example from the WorldTree dataset of how the explanation changes after each iteration. Figures 36, 37, 38, 39, 40, 41, 42 and 43 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 8 and Figures 44, 46, 47 and 48, as well as Table 9 with Figures 49, 50, 51, and 52, provide two more examples.
A.5.4 Unrefined Example
Table 10 and Table 11 shows an example from the WorldTree dataset that does not refine within 10 iterations and is not caused by a syntax error. The figures that follow show the detailed Isabelle/HOL theory code of the related iterations.
Figure 11: Average Progressed Explanations against Number of Planned Explanations in Refined and Unrefined e-SNLI, QASC and WorldTree Dataset
Figure 12: Average Progressed Explanations against Number of Planned Explanations for Refined, Unrefined, and Combined Across All Datasets
Figure 13: Average Processed Proof Steps against Total Suggested Proof Steps for Refined, Unrefined, and Combined Across All Datasets
1
2
3
Input : Premise $p$, Explanation $E$, Hypothesis $h$, Isabelle//HOL server $isabelle$, Autoformalisation model $m_{a}$, Isabelle syntax refinement model $m_{sr}$, Rough inference model $m_{ri}$, Proof step build model $m_{pr}$, Facts filter model $m_{f}$, Explanation refinement model $m_{e}$
Output : Updated Explanation $E$
4 $valid$ $\leftarrow false$
5 isabelle_theory [$\leftarrow$[\ ]← [ ]
6 $iterations\leftarrow 0$
7 $max\_iterations\leftarrow 11$
8 has_syntax_error $\leftarrow$ $false$
9 while not valid and $iterations<max\_iterations$ do
10       session_id $\leftarrow$ session_build($HOL,isabelle$)
11       $isabelle$.start(session_id)
12       isabelle_theory $\leftarrow$ transfer_to_symbolic($p,E,h$, $m_{a}$)
13       messages, error_content, error_code $\leftarrow$ $isabelle$.check(isabelle_theory)
14       if syntax_errors in messages then
15             has_syntax_error $\leftarrow true$
16             $it\leftarrow 0$
17             while has_syntax_error and $it<3$ do
18                   isabelle_theory = refine_syntax(messages, error_content, error_code, isabelle_theory, $m_{sr}$)
19                   messages, error_content, error_code $\leftarrow$ $isabelle$.check(isabelle_theory)
20                   if syntax_errors in messages then
21                         has_syntax_error $\leftarrow true$
22                         $it\leftarrow it+1$
23
24                  else
25                         break
26                   end if
27
28             end while
29
30       end if
31      rough_inference $\leftarrow$ make_rough_inference($p,E,h,m_{ri}$)
32       proof_steps $\leftarrow$ build_proof(rough_inference, $m_{pr}$)
33       isabelle_theory $\leftarrow$ isabelle_theory + proof_steps
34       messages, error_content, error_code $\leftarrow$ $isabelle$.check(isabelle_theory)
35       if messages is not empty then
36             message $\leftarrow$ messages[0]
37             $E$leftarrow filter($E$, rough_inference, proof_steps, $m_{f}$)
38             $E$leftarrow refine_explanation(message, error_content, error_code, rough_inference, proof_steps, $p,E,H,m_{e}$)
39
40      else
41             $valid\leftarrow true$
42             break
43       end if
44      $iterations\leftarrow iterations+1$
45       $isabelle$.shutdown()
46
47 end while
return $E$
Algorithm 1 Explanation-Refiner
Figure 14: Prompts for detecting event-related words in the given sentences
Figure 15: Prompts for converting natural language sentences into logical form representations
Figure 16: Prompts for converting logical form into Isabelle/HOL code format for building the axioms and type declaration
Figure 17: Prompts for building the theorem code part of the Isabelle/HOL theory
Figure 18: Prompts for how to refine the identified syntax errors in the constructed code
Figure 19: Prompts for how to make a step-by-step preliminary inference strategy
Figure 20: Prompts for how to build a proof for Isabelle/HOL proof assistant
Figure 21: Prompts for how to refine the explanatory sentences
Dataset Sentences Explanation Iteration Validity
e-SNLI Premise: A woman in black framed glasses peruses a photo album while sitting in a red wicker chair.
Hypothesis: There is a lady with a book.
The lady is looking through a photo album which is a type of book. 0 Invalid
e-SNLI Premise: A woman in black framed glasses peruses a photo album while sitting in a red wicker chair.
Hypothesis: There is a lady with a book.
A woman can be referred to as a lady. A photo album is a type of book. 1 Invalid
e-SNLI Premise: A woman in black framed glasses peruses a photo album while sitting in a red wicker chair.
Hypothesis: There is a lady with a book.
A woman can be referred to as a lady. A photo album is a type of book. If a woman is perusing a photo album, then the woman is with a book. 2 Valid
Table 1: An example of how the explanation sentences in e-SNLI can be refined with Explanation-Refiner
Figure 22: The Isabelle theory code for table 1 iteration 0
Figure 23: The Isabelle theory code for table 1 iteration 1
Figure 24: The Isabelle theory code for table 1 iteration 2
Dataset Sentences Explanation Iteration Validity
e-SNLI Premise: A couple being romantic under the sunset.
Hypothesis: People are enjoying sundown.
A couple consists of people. Sunset is sundown. 0 Invalid
e-SNLI Premise: A couple being romantic under the sunset.
Hypothesis: People are enjoying sundown.
A couple consists of people. Sunset is sundown. Being romantic under the sunset implies enjoying the sundown. 1 Valid
Table 2: An example of how the explanation sentences in e-SNLI can be refined with Explanation-Refiner
Figure 25: The Isabelle theory code for table 2 iteration 0
Figure 26: The Isabelle theory code for table 2 iteration 1
Dataset Sentences Explanation Iteration Validity
e-SNLI Premise: A male bartender dressed in all black with his sleeves rolled up to elbow height making a drink in a martini glass.
Hypothesis: A person in black
A bartender, who is a person, is wearing black. 0 Invalid
e-SNLI Premise: A male bartender dressed in all black with his sleeves rolled up to elbow height making a drink in a martini glass.
Hypothesis: A person in black
A bartender is a person. If a person is wearing black, then the person is in black. 1 Invalid
e-SNLI Premise: A male bartender dressed in all black with his sleeves rolled up to elbow height making a drink in a martini glass.
Hypothesis: A person in black
A bartender is a person. If a person is dressed in black, then the person is in black. 2 Valid
Table 3: An example of how the explanation sentences in e-SNLI can be refined with Explanation-Refiner
Figure 27: The Isabelle theory code for table 3 iteration 0
Figure 28: The Isabelle theory code for table 3 iteration 1
Figure 29: The Isabelle theory code for table 3 iteration 2
Dataset Sentences Explanation Iteration Validity
QASC Hypothesis: Meiosis is the type of cell division that produces haploid cells. Meiosis is the type of cell division that produces gametes. Gametes are haploid egg or sperm that fuse to form a zygote. 0 Invalid
QASC Hypothesis: Meiosis is the type of cell division that produces haploid cells. Meiosis is the type of cell division that produces gametes. Gametes are haploid cells. Haploid cells are either egg or sperm that fuse to form a zygote. 1 Valid
Table 4: An example of how the explanation sentences in QASC can be refined with Explanation-Refiner
Figure 30: The Isabelle theory code for table 4 iteration 0
Figure 31: The Isabelle theory code for table 4 iteration 1
Dataset Sentences Explanation Iteration Validity
QASC Hypothesis: Antigens are found on cancer cells and the cells of transplanted kidneys. Antigens are found on cancer cells and the cells of transplanted organs. Examples of animal organs are kidneys and hearts. 0 Invalid
QASC Hypothesis: Antigens are found on cancer cells and the cells of transplanted kidneys. Antigens are found on cancer cells and the cells of transplanted organs. Kidneys are examples of animal organs. Transplanted organs include kidneys. 1 Valid
Table 5: An example of how the explanation sentences in QASC can be refined with Explanation-Refiner
Figure 32: The Isabelle theory code for table 5 iteration 0
Figure 33: The Isabelle theory code for table 5 iteration 1
Dataset Sentences Explanation Iteration Validity
QASC Hypothesis: Azathioprine interferes with the assembly of proteins. DNA contains genetic instructions for proteins, and RNA helps assemble the proteins. Azathioprine interferes with DNA and RNA synthesis 0 Invalid
QASC Hypothesis: Azathioprine interferes with the assembly of proteins. DNA contains genetic instructions for proteins, and RNA, synthesized from DNA, helps assemble the proteins. The assembly of proteins is dependent on the synthesis of DNA and RNA. Azathioprine interferes with DNA and RNA synthesis. 1 Valid
Table 6: An example of how the explanation sentences in QASC can be refined with Explanation-Refiner
Figure 34: The Isabelle theory code for table 6 iteration 0
Figure 35: The Isabelle theory code for table 6 iteration 1
Dataset Sentences Explanation Iteration Validity
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. when an electrical conductor is plugged into an outlet , a circuit is completed. a complete electrical circuit is a source of electrical energy. electricity means electrical energy. safe is the opposite of harm; danger. plugging many appliances into one outlet is dangerous. to be in danger means to be around; to be near something dangerous. electrocution causes harm to an organism. if electricity flows through; is transferred through the body of an animal then that animal is electrocuted. an animal is a kind of organism. 0 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. 1 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. An event is considered safe around electricity if it is safe. 2 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. An event is considered not safe around electricity if it is not safe. 3 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. If an event is not safe, it is considered not safe around electricity. 4 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. If an event related to electricity is not safe, it is considered not safe around electricity. 5 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. Plugging many appliances into one outlet is an event related to electricity. If an event related to electricity is not safe, it is considered not safe around electricity. 6 Invalid
WorldTree Hypothesis: Plugging many appliances into one outlet is not a way to stay safe around electricity. if an event is dangerous, it is not safe. plugging many appliances into one outlet is dangerous. Plugging many appliances into one outlet is an event related to electricity. If an event related to electricity is not safe, it is considered not safe around electricity. If an event is not safe around electricity, it is not a way to stay safe around electricity. 7 Valid
Table 7: An example of how the explanation sentences in WorldTree can be refined with Explanation-Refiner
Figure 36: The Isabelle theory code for table 7 iteration 0
Figure 37: The Isabelle theory code for table 7 iteration 1
Figure 38: The Isabelle theory code for table 7 iteration 2
Figure 39: The Isabelle theory code for table 7 iteration 3
Figure 40: The Isabelle theory code for table 7 iteration 4
Figure 41: The Isabelle theory code for table 7 iteration 5
Figure 42: The Isabelle theory code for table 7 iteration 6
Figure 43: The Isabelle theory code for table 7 iteration 7
Dataset Sentences Explanation Iteration Validity
WorldTree Hypothesis: Sound energy can travel by vibrating particles of air. sound is a kind of energy. sound is a kind of air vibration. sound can travel through air. form means kind. 0 Invalid
WorldTree Hypothesis: Sound energy can travel by vibrating particles of air. Sound energy is a kind of energy. Sound is a kind of air vibration. Sound can travel through air. Form means kind. 1 Invalid
WorldTree Hypothesis: Sound energy can travel by vibrating particles of air. Sound energy is the energy that sound carries. Sound is a kind of air vibration. Sound can travel through air by vibrating particles of air. 2 Invalid
WorldTree Hypothesis: Sound energy can travel by vibrating particles of air. Sound energy is the energy that sound carries. Sound energy is a kind of air vibration. Sound can travel through air by vibrating particles of air. 3 Invalid
WorldTree Hypothesis: Sound energy can travel by vibrating particles of air. Sound energy is the energy that sound carries. Sound energy can cause air vibration. Sound can travel through air by vibrating particles of air. 4 Valid
Table 8: An example of how the explanation sentences in WorldTree can be refined with Explanation-Refiner
Figure 44: The Isabelle theory code for table 8 iteration 0
Figure 45: The Isabelle theory code for table 8 iteration 1
Figure 46: The Isabelle theory code for table 8 iteration 2
Figure 47: The Isabelle theory code for table 8 iteration 3
Figure 48: The Isabelle theory code for table 8 iteration 4
Dataset Sentences Explanation Iteration Validity
WorldTree Premise: most birds sit on their eggs until they hatch.
Hypothesis: The reason why birds sit on their eggs is to keep the eggs warm.
an egg requires warmth to survive; hatch; develop. sitting on something provides warmth to that something. providing warmth is synonymous with keeping warm. form means kind. 0 Invalid
WorldTree Premise: most birds sit on their eggs until they hatch.
Hypothesis: The reason why birds sit on their eggs is to keep the eggs warm.
Eggs require warmth to survive, hatch, and develop. Sitting on something provides warmth to that something. Providing warmth is synonymous with keeping warm. 1 Invalid
WorldTree Premise: most birds sit on their eggs until they hatch.
Hypothesis: The reason why birds sit on their eggs is to keep the eggs warm.
Birds sitting on their eggs provides warmth to the eggs. Providing warmth is synonymous with keeping warm. 2 Invalid
WorldTree Premise: most birds sit on their eggs until they hatch.
Hypothesis: The reason why birds sit on their eggs is to keep the eggs warm.
Birds sit on their eggs to provide warmth to the eggs. Providing warmth is synonymous with keeping warm. 3 Valid
Table 9: An example of how the explanation sentences in WorldTree can be refined with Explanation-Refiner
Figure 49: The Isabelle theory code for table 9 iteration 0
Figure 50: The Isabelle theory code for table 9 iteration 1
Figure 51: The Isabelle theory code for table 9 iteration 2
Figure 52: The Isabelle theory code for table 9 iteration 3
Dataset Sentences Explanation Iteration Validity
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
reflecting light; sound is the opposite of absorbing light; sound. if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something is in the sunlight then that object; that something will absorb solar energy. shiny things; objects reflect light. keeping cool means not increasing temperature. a car is a kind of object. 0 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something is in the sunlight then that object; that something will absorb solar energy. shiny things; objects reflect light. keeping cool means not increasing temperature. a parked car is a kind of object. 1 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something is in the sunlight and it does not reflect light, then that object; that something will absorb solar energy. shiny things; objects reflect light. keeping cool means not increasing temperature. a parked car is a kind of object. A shiny aluminum screen is a kind of shiny object. 2 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something is in the sunlight and it reflects light, then that object; that something will not absorb solar energy. shiny things; objects reflect light. keeping cool means not increasing temperature. A shiny aluminum screen is a kind of shiny object. 3 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something is in the sunlight and it reflects light, then that object; that something will not absorb solar energy. A shiny aluminum screen reflects light. keeping cool means not increasing temperature. A shiny aluminum screen is a kind of shiny object. 4 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something reflects light, then that object; that something will not absorb solar energy. A shiny aluminum screen reflects light. keeping cool means not increasing temperature. 5 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
if an object; a substance absorbs solar energy then that object; that substance will increase in temperature. if an object; something reflects light, then that object; that something will not absorb solar energy. A shiny aluminum screen reflects sunlight. keeping cool means not increasing temperature. 6 Invalid
Table 10: An example of an explanation in WorldTree that does not refine within 10 iterations using Explanation-Refiner
Dataset Sentences Explanation Iteration Validity
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
If a shiny aluminum screen absorbs solar energy then that shiny aluminum screen will increase in temperature. If a shiny aluminum screen reflects light, then that shiny aluminum screen will not absorb solar energy. A shiny aluminum screen reflects sunlight. Keeping cool means not increasing temperature. 7 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
If a shiny aluminum screen absorbs solar energy then that shiny aluminum screen will increase in temperature. If a shiny aluminum screen reflects sunlight, then that shiny aluminum screen will not absorb solar energy. A shiny aluminum screen reflects sunlight. Keeping cool means not increasing temperature. 8 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
If a shiny aluminum screen absorbs solar energy then that shiny aluminum screen will increase in temperature. A shiny aluminum screen that reflects sunlight does not absorb solar energy. A shiny aluminum screen reflects sunlight. Keeping cool means not increasing temperature. 9 Invalid
WorldTree Premise: A shiny aluminum screen can be placed on the windshield of a parked car.
Hypothesis: The screen helps to keep the car cool because it reflects the sunlight.
If a shiny aluminum screen absorbs solar energy then that shiny aluminum screen will increase in temperature. A shiny aluminum screen that reflects sunlight does not absorb any energy. A shiny aluminum screen reflects sunlight. Keeping cool means not increasing temperature. 10 Invalid
Table 11: An example of an explanation in WorldTree that does not refine within 10 iterations using Explanation-Refiner
Figure 53: The Isabelle theory code for table 10 iteration 0
Figure 54: The Isabelle theory code for table 10 iteration 1
Figure 55: The Isabelle theory code for table 10 iteration 2
Figure 56: The Isabelle theory code for table 10 iteration 3
Figure 57: The Isabelle theory code for table 10 iteration 4
Figure 58: The Isabelle theory code for table 10 iteration 5
Figure 59: The Isabelle theory code for table 10 iteration 6
Figure 60: The Isabelle theory code for table 11 iteration 7
Figure 61: The Isabelle theory code for table 11 iteration 8
Figure 62: The Isabelle theory code for table 11 iteration 9
Figure 63: The Isabelle theory code for table 11 iteration 10